OpenStudy (anonymous):

For this equation: 2x + 5y = 10, does x = 5 and y = 2?

OpenStudy (anonymous):

just substitute values & check; 2*5 +5*2 = ... right or wrong?

OpenStudy (anonymous):

x = 10, y = -2

OpenStudy (anonymous):

x-y = (-2/5)x-2 y = (-2/5)x + 2 y = -2x + 10 or y = (2/5)x -2 i don know what u trying to do but this what i got

OpenStudy (anonymous):

I'm sorry. I meant to put in, for finding the horizontal intercept and vertical intercept of the equation.

OpenStudy (anonymous):

2(5)+5(2)=10 10+10=10? No =20

OpenStudy (anonymous):

I don't get how to find the horizontal and vertical intercepts then.. I substituted y for 0 and vice versa for x.

OpenStudy (anonymous):

its correct, they are the intercepts

OpenStudy (anonymous):

x = 5 and y = 2?